# GAMA: Generative Agents for Multi-Agent Autoformalization

### v2

@@ 4.1 Parameters
Table 2: Common experimental parameters.
@@
match maker S S RR S
Table 3: Experiment-specific parameters. S stands for a mode in which an agent plays against its clone with an opposite strategy, while RR stands for round robin.
@@ 4.2 Autoformalization of Interaction Descriptions (Numerical Payoffs)
We used a dataset of 55 natural language game-theoretic scenarios with numerical payoffs to evaluate the syntactic and semantic accuracy of generated logic programs. The dataset included 5 scenarios that employed common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games introduced in Section 2.1. The remaining 50 scenarios—10 for each game—used alternative descriptions that avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma). For each scenario, five agents autoformalized game rules and then played tournament against their clones, generating all four possible pairs of moves in four rounds. Each agent had a maximum of five attempts to produce syntactically correct code; agents failing after five attempts were labelled as syntactically incorrect. Post-tournament, each agent’s individual and total payoffs were compared to the target, with matches indicating semantic correctness.
@@ 4.3 Autoformalization of Interaction Descriptions (Non-numerical Payoffs)
This experiment was analogous to the previous one, but it utilized interaction scenarios without numerical payoffs (e.g., ”If both remain silent, they get minor sentences,” to describe the outcome of mutual cooperation in PD). Therefore, the autoformalizing LLM had to infer the appropriate payoffs correctly. Since it was not possible to provide a ground truth matrix in this case, we validated the semantic correctness by performing constraint checks (described in Sect. 3.2) to determine whether the payoff matrix satisfied the constraints for the ground truth game type (Sect. 2.1).
@@ 4.4 Axelrod’s Tournament
Table 4: Strategies utilized in the evaluation.
To demonstrate the framework’s usefulness in analysing interaction scenarios, we selected five agents created during the first experiment, one for each of the five games. For each game, we conducted a tournament modelled after Axelrod’s tournament Axelrod and Hamilton (1981). In this setup, six copies of each agent were created and assigned one of six strategies. Each agent then played against every other agent, including itself. The strategies used are listed in Table 4.
The main metric for evaluating strategy performance was the total payoff achieved by an agent using a given strategy in each game. To allow for cross-game comparisons of strategy effectiveness, total payoffs were normalized using:
$norm\_payoff=\frac{total\_payoff-min\_total\_payoff}{{max\_total\_payoff-min\_total\_payoff}},$
$Total\_payoff$ refers to the sum of the payoffs an agent collects across all rounds and $min\_total\_payoff$ and $max\_total\_payoff$ represent the minimum and maximum total payoffs achievable, respectively.
@@ 4.5 Autoformalization of Strategies
We then evaluated the semantic correctness of the autoformalized strategies. To achieve this, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as an example. For the other five strategies, the autoformalization module was provided with their natural language descriptions, similar to those shown in Table 4.
Each autoformalized strategy was then assigned to an agent with predefined Prisoner’s Dilemma game rules. This agent played four rounds against a clone utilizing the anti-tit-for-tat strategy. The agent’s total payoff was compared to a target payoff specific to each strategy to assess correctness. For the random strategy, correctness was also manually validated. This tournament setup was repeated five times.
@@ 5.1 Autoformalization of Interaction Descriptions
1 2 3 4 5
@@
Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
GPT-4o 97.64% 2.02% 0.34% 0.00% 0.00%
Table 5: Distribution of attempts at creating syntactically correct code for each model.
Table 5 shows the distribution of attempts needed to generate syntactically correct code for both models. The type of description (numeric vs. non-numeric) had no influence on syntactic correctness of the generated code. Claude 3.5 Sonnet achieved syntactically correct code in 100% of cases and GPT-4o in nearly 100%. The few errors observed involved non-Prolog comment characters and non-escaped special characters in strings, all of which were successfully corrected on the second attempt. This demonstrates that providing a specific line that generated the error in the feedback prompt significantly enhances an LLM’s ability to self-correct compared to merely providing an error message Mensfelt et al. (2024). The only case where GPT-4o was unable to correct a syntactic error involved the error message “Unexpected end of file“ which indicated the end of the file as the error’s source. However, the actual line generating the error was not provided in the feedback prompt.
Syntactic Runtime Semantic
@@
Numeric Payoffs
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 89.1% 96.4% 68.5% 96.4% 87.3%
Semantic 89.1% 96.4% 68.5% 89.1% 87.3%
Non-Numeric Payoffs
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 82.1% 85.7% 74.0% 70.7% 83.1%
Semantic 82.1% 77.8% 34.2% 64.0% 83.1%
@@
BS HD MP PD SH
Syntactic 98.3% 100.0% 100.0% 100.0% 100.0%
Runtime 83.1% 89.1% 96.4% 89.1% 90.9%
Semantic 79.7% 89.1% 90.9% 85.5% 83.6%
@@
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 93.2% 88.1% 87.3% 96.5% 91.5%
Semantic 93.2% 67.8% 30.2% 73.7% 76.3%
@@
The semantic correctness was lower than runtime correctness for both models and description types (Table 6). This means that even though an agent was able to select a move during runtime and update its state, the semantics of the program were not correct. For descriptions with numerical payoffs, this implies that the payoff matrix does not contain the required payoffs. For descriptions without numerical payoffs, it indicates that the payoff matrix does not satisfy the constraints of the required game. Tables 7 and 8 present the syntactic, runtime, and semantic correctness for different games.
The runtime and semantic correctness varied depending on the model and the game. Notably, Matching Pennies consistently resulted in low semantic correctness for non-numerical descriptions, with 34.2% correctness for Claude 3.5 Sonnet and 30.2% correctness for GPT-4o. This may be attributed to some descriptions specifying which player wins if the choices match or mismatch, without explicitly stating that the other player will lose—a constraint of the game. Extending the autoformalization feedback loop to include semantic evaluation has the potential to significantly improve correctness.
@@ 5.2 Axelrod’s Tournament
Figure 2: Normalized total payoff for each strategy by game.
Figure 2 presents the normalized total payoffs for each strategy across all games. Overall, the best-response strategy proved to be the most effective on average, followed closely by tit-for-tat. Best-response was most effective in all games except the Battle of the Sexes and the Prisoner’s Dilemma. In the Prisoner’s Dilemma, best-response achieved a lower payoff than the anti-default-move strategy (analogous to “defect”), as it initiated with the “cooperate” move. The results demonstrate that agents with autoformalized rules can effectively simulate and assess the performance of various strategies within different modelled scenarios.
@@ 5.3 Autoformalization of Strategies
Strategy Claude 3.5 Sonnet GPT-4o
anti-default-move 100.00% 100.00%
anti-tit-for-tat 100.00% 100.00%
best-response 80.00% 100.00%
default-move 100.00% 100.00%
random 80.00% 100.00%
Table 9: Percentage of semantic correctness for each strategy.
Table 9 presents the percentage of correctly autoformalized strategies. GPT-4o achieved 100% correctness in autoformalizing all five strategies. Claude 3.5 Sonnet produced one incorrect strategy for best-response and one for random, which are the strategies that deviate the most from the tit-for-tat strategy presented as an example in the prompt. One error occurred during runtime, and another during semantic validation. Overall, the high correctness demonstrates that the GAMA framework has the potential to effectively autoformalize both game rules and gameplay strategies.
6 Related Work
Originally proposed for agents in game scenarios, the Game Description Language (GDL) offers a promising approach for formalising complex strategic interactions in multi-agent environments by providing a declarative framework to define game rules, legal moves, and outcome conditions Genesereth et al. (2005). Agents that employ GDL are often used to systematically explore game states and strategic options, supporting rigorous analysis across both cooperative and competitive games, as suggested by previous approaches (e.g., see  Schiffel and Thielscher (2009)). However, these approaches typically rely on human-driven formalisation of specific games, a limitation addressed by the LLM-augmented agents and the GDL representation used here, which enable autoformalization and enhance flexibility.
There is a growing body of research exploring the use of LLMs in game theory Lorè and Heydari (2024), systematically examining their potential as player agents in game-theoretic scenarios Fan et al. (2024), utilising their inference abilities to make decisions in games with imperfect information Guo et al. (2023), comparing their behaviour with human participants Fontana et al. (2024), and deploying them in simulation frameworks Mao et al. (2024). While we share with these works an interest in the broad capabilities and value of LLMs in game theoretic settings, our focus diverges by concentrating on how to use LLMs to generate GDL-like executable game specifications. These formal specifications enable agents to rigorously analyse game situations and reason about possible outcomes, providing a structured and robust foundation for decision-making within game-theoretic frameworks. Our work builds on previous research discussed in Mensfelt et al. (2024), significantly extending it by embedding the autoformalization module within an agent model, incorporating the autoformalization of player strategies, implementing a distinct correctness validation process for the autoformalization through tournament play, and conducting experiments to validate the results.
Autoformalization Wu et al. (2022), has shown promising results in translating natural language into mathematical formalisms, enabling the integration of large language models with formal tools for solving reasoning tasks Wu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022). This approach has also been successfully extended to creating logical representations from natural language descriptions. For instance, Yang et al. (2023) applied GPT-3 with few-shot prompting to translate natural language sentences into formal representations, which were then inputted into answer set programs. While effective on NLP benchmarks, this method faced occasional translation errors. To further enhance performance, Pan et al. (2023) combined in-context learning with self-refinement based on solver evaluations, surpassing standard LLM and chain-of-thought prompting approaches on logical reasoning benchmarks. In the context of temporal logic, Cosler et al. (2023) addressed the problem of translation errors with an interactive system allowing users to refine sub-translations. Fine-tuning LLMs has also proven effective for boosting accuracy in translations to temporal logics Chen et al. (2023) and deductive datasets Feng et al. (2023).
@@ 7 Conclusions and Future Work
We have presented a framework for the autoformalization of interaction scenarios, specifically those modelled by $2\times 2$ simultaneous-move games. The framework enables the autoformalization of strategies for these games and the simulation of tournaments based on the autoformalized rules. The generated programs undergo three levels of validation: syntactic validation by consulting a solver, functional testing through gameplay, and semantic validation by comparing achieved payoffs with target outcomes or by checking the game constraints. Evaluation of the framework demonstrated high syntactic and semantic accuracy in the code generated by agents across various interaction scenarios.
Future work will focus on extending the framework to model games beyond $2\times 2$ simultaneous-move scenarios. Our previous work Mensfelt et al. (2024) demonstrated that using our sample solver for the Prisoner’s Dilemma we could successfully generate code for structurally similar games, such as rock-paper-scissors (with a different number of moves) and sequential versions of the PD. However, achieving successful generalization to a broader class of games will require assembling a more diverse set of in-context learning examples and developing methods to retrieve the most relevant examples for autoformalization in specific scenarios. Another direction for future work involves extending the autoformalization feedback loop to include runtime and semantic validation checks, further improving the correctness of the generated programs.
These extensions will provide a robust framework for generating reasoning modules for decision-making agents in a wider class of games and game-like interactions.
Acknowledgments

### v3

@@ 4.1 Parameters
Table 3: Experiment-specific parameters. C stands for a mode in which an agent plays against its clone with an opposite strategy, while RR stands for round robin.
@@
Match maker C C RR C
@@
4.2 Dataset
We used a dataset of 110 natural language game-theoretic scenarios, with and without numerical payoffs, based on an improved version of the dataset introduced in [23] (available in the repository). For both variants, the dataset included 5 scenarios per game employing common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games described in Section 2.1. The remaining 50 scenarios—10 per game—used alternative descriptions that deliberately avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma).
@@ 4.3 Autoformalization of Interaction Descriptions (Numerical Payoffs)
For each scenario, five agents autoformalized the game rules and subsequently played tournaments against clones of themselves, generating all four possible pairs of moves across four rounds. Each agent was allowed up to five attempts to produce syntactically correct code; agents that failed after five attempts were labeled as syntactically incorrect. Following the tournament, each agent’s individual and total payoffs were compared to the target values, with exact matches indicating semantic correctness.
@@ 4.4 Autoformalization of Interaction Descriptions (Non-Numerical Payoffs)
This experiment followed the same procedure as the numerical payoffs evaluation but used interaction scenarios without explicit numerical payoffs (e.g., “If both remain silent, they receive minor sentences” to describe mutual cooperation in the Prisoner’s Dilemma). In these cases, the autoformalizing LLM was required to infer appropriate payoffs from the textual descriptions. As a ground truth payoff matrix could not be provided, semantic correctness was instead assessed through constraint checking (as described in Section 3.5), verifying whether the generated payoff matrix satisfied the properties required for the corresponding game type (Section 2.1).
@@ 4.5 Axelrod’s Tournament
To demonstrate the framework’s applicability for analysing interaction scenarios, we selected five agents generated during the first experiment, one representing each of the five canonical games. For each game, we conducted a tournament inspired by Axelrod’s seminal tournament [4]. In this setup, six copies of each agent were created and assigned one of six strategies. Each agent competed against every other agent, including itself. The strategies are listed in Table 4.
The primary metric for evaluating strategy performance was the total payoff achieved by an agent using a particular strategy within each game. To enable cross-game comparisons of strategy effectiveness, total payoffs were normalized using the following formula:
$\mathrm{NormPayoff}=\frac{\mathrm{TotalPayoff}-\mathrm{MinTotalPayoff}}{\mathrm{MaxTotalPayoff}-\mathrm{MinTotalPayoff}}$
where TotalPayoff denotes the sum of payoffs accumulated by an agent across all rounds, and MinTotalPayoff and MaxTotalPayoff represent the minimum and maximum total payoffs achievable in that game, respectively.
@@ 4.6 Autoformalization of Strategies
For the autoformalization of strategies, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as an example for one-shot learning. The autoformalization module then received natural language descriptions of the remaining five strategies, similar to those shown in Table 4, to generate their representations.
Each autoformalized strategy was assigned to an agent equipped with predefined Prisoner’s Dilemma rules. The agent then played four rounds against a clone using the anti-tit-for-tat strategy. As in the game autoformalization experiments, the agent’s total payoff was compared against a target payoff specific to each strategy to assess semantic correctness. For the random strategy, correctness was additionally verified through manual inspection.
@@ 5.1 Autoformalization of Interaction Descriptions
Table 5: Percentage of number of attempts at creating syntactically correct code for each model.
1 (%) 2 (%) 3 (%) 4 (%) 5 (%)
@@
Claude 3.5 S $100.00$ $0.00$ $0.00$ $0.00$ $0.00$
GPT-4o $97.64$ $2.02$ $0.34$ $0.00$ $0.00$
Table 5 presents the distribution of attempts required to generate syntactically correct code for both models. Claude 3.5 Sonnet achieved syntactically correct code on the first attempt in 100% of cases, while GPT-4o did so in 97% of cases.
The few errors observed in GPT-4o primarily involved non-Prolog comment characters or unescaped special characters in strings, most of which were corrected on the second attempt. These results demonstrate that providing the specific Prolog line that triggered the error in the feedback prompt significantly improves the LLM’s ability to self-correct, compared to providing only an error message [23].
The only instance where GPT-4o failed to correct a syntactic error involved an “Unexpected end of file” message. In this case, the feedback prompt did not include the problematic line itself, but only the generic error message, limiting the model’s ability to identify and fix the issue.
Syntactic (%) Runtime (%) Semantic (%)
@@
BS (%) HD (%) MP (%) PD (%) SH (%)
Syntactic $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
Runtime $89.1$ $96.4$ $68.5$ $96.4$ $87.3$
Semantic $89.1$ $96.4$ $68.5$ $89.1$ $87.3$
@@
BS (%) HD (%) MP (%) PD (%) SH (%)
Syntactic $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
Runtime $82.1$ $85.7$ $74.0$ $70.7$ $83.1$
Semantic $82.1$ $77.8$ $34.2$ $64.0$ $83.1$
@@
Numeric Payoffs
BS (%) HD (%) MP (%) PD (%) SH (%)
Syntactic $98.3$ $100.0$ $100.0$ $100.0$ $100.0$
Runtime $83.1$ $89.1$ $96.4$ $89.1$ $90.9$
Semantic $79.7$ $89.1$ $90.9$ $85.5$ $83.6$
Non-Numeric Payoffs (%)
BS (%) HD (%) MP (%) PD (%) SH (%)
Syntactic $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
Runtime $93.2$ $88.1$ $87.3$ $96.5$ $91.5$
Semantic $93.2$ $67.8$ $30.2$ $73.7$ $76.3$
Semantic correctness was lower than runtime correctness for both models and description types (Table 6). This indicates that, although an agent was able to select moves and update its state during execution, the underlying semantics of the program were incorrect. For descriptions with numerical payoffs, semantic errors imply that the generated payoff matrix did not contain the required values. For descriptions without numerical payoffs, they indicate that the payoff matrix failed to satisfy the constraints associated with the intended game. Detailed syntactic, runtime, and semantic correctness results for different games are shown in Tables 7 and 8.
Runtime and semantic correctness varied across models and games. Notably, Matching Pennies consistently exhibited low semantic correctness for non-numerical descriptions, with only 34.2% correctness for Claude 3.5 Sonnet and 30.2% for GPT-4o. This may be attributed to certain descriptions specifying only which player wins when choices match or mismatch, without explicitly stating that the other player loses—a crucial constraint of the game. Incorporating semantic evaluation into the autoformalization feedback loop may substantially improve correctness by addressing such issues.
@@ 5.2 Axelrod’s Tournament
Figure 2: Normalized total payoff for each strategy by game. Rows are ordered by the strategies with the highest average total payoff, and columns are arranged similarly based on the games. The total payoff for the winning strategy in each game is highlighted in white.
Figure 2 presents the normalized total payoffs for each strategy across all games. Overall, the best-response strategy proved the most effective on average, followed closely by tit-for-tat. Best-response achieved the highest scores in Stag Hunt and Matching Pennies (tied with tit-for-tat in the latter). However, in the Prisoner’s Dilemma, it earned a lower payoff than the anti-default-move strategy (analogous to “Defect”), as it opened with a “Cooperate” move.
These results demonstrate that agents equipped with autoformalized rules can successfully simulate strategic interactions and assess the relative performance of various strategies across different modelled scenarios.
@@ 5.3 Autoformalization of Strategies
Table 9: Percentage of semantic correctness for each autoformalized strategy.
Strategy Claude 3.5 Sonnet (%) GPT-4o (%)
anti-default-move $100.00$ $100.00$
anti-tit-for-tat $100.00$ $100.00$
best-response $80.00$ $100.00$
default-move $100.00$ $100.00$
random $80.00$ $100.00$
Table 9 presents the percentage of correctly autoformalized strategies. GPT-4o achieved 100% correctness across all five strategies. Claude 3.5 Sonnet produced one incorrect autoformalization each for best-response and random, which are the most divergent from the tit-for-tat strategy provided as an example in the prompt. Of these errors, one occurred during runtime and the other during semantic validation.
Overall, the high levels of correctness demonstrate that the GAMA framework can effectively autoformalize both game rules and gameplay strategies, including those with varying levels of similarity to the examples provided.
@@ 6 Conclusions and Future Work
We have proposed a framework for the autoformalization of interaction scenarios, focusing on those modelled as $2\times 2$ simultaneous-move games. The framework supports autoformalization of both game rules and strategies, and enables simulation of tournaments based on the generated formal representations. Programs produced by the framework undergo three levels of validation: syntactic validation through solver consultation, runtime testing via gameplay, and semantic validation based on target payoffs or game-specific constraints. Experimental results demonstrated high syntactic and semantic accuracy across diverse interaction scenarios.
Future work will extend the framework beyond $2\times 2$ simultaneous-move games. Our previous work [23] showed that generalization to structurally similar games, such as rock-paper-scissors and sequential versions of the Prisoner’s Dilemma, could be addressed using the same solver. Generalization to a broader class of games, however, will require assembling a more diverse collection of in-context learning examples and developing retrieval mechanisms to select relevant examples for autoformalization in specific scenarios. Another promising direction involves expanding the feedback loop to incorporate runtime and semantic validation, further enhancing the correctness and robustness of the generated programs.
These extensions aim to establish a more general and reliable framework for generating reasoning modules, enabling decision-making agents to operate effectively across a wider range of games and interaction scenarios.
{ack}